[Hallc_running] [Revised Logentry] Hall C status and plan
ciprian at jlab.org
ciprian at jlab.org
Fri Oct 8 17:55:02 EDT 2021
* Beam delivery has been very good, with the occasional separator problem.
* For a short period of time last night the accelerator could not establish an energy lock so they monitored and manually adjusted (https://logbooks.jlab.org/entry/3919887).
* We've had a couple of delays being angle to enter the hall since the last update relating to the ozone monitor (see details about accesses below) which are hoped to be remedied by a new monitor that will be installed soon. In the meantime MCC has retained control over the handheld monitor that will allow them to test the ozone level before making an quick access.
* An air hose was split on a beamline valve close to the target chamber and technical staff managed to cut it and splice it together. On the next long down we should plan to replace it with a hard line.
* Hall A setup will start Monday 9AM; we should expect at least 4h of no-to-little beam. Currently we plan to remain closed up for the duration of the setup.
* After about 14h of downtime we have now started taking beam again (https://logbooks.jlab.org/entry/3920347). At about 2AM the shift crew noticed that all the interlocks tripped for the SHMS (https://logbooks.jlab.org/entry/3920023). This was due to the loss of power to the cryo controlers in the SHMS. The UPS that fed all the controllers died (see att figure). During the access Ellen tried to reboot it by flipping the breaker but was not successful. It indicates that there is power going in but we are not able to control it or even bring up the display screen. The target on-call (thanks Silviu) was called to see if we can prevent the target from warming up or damaging it (https://logbooks.jlab.org/entry/3920042). At his request we contacted cryo (https://logbooks.jlab.org/entry/3920047) to let them know what was going on and see if there is anything we could do. Because the cryo controls were unreachable nothing could be done before accesing the hall. Luckly, hall C was th!
e only one using serious amount of cooling so we did not end up tripping the ESR. The HMS tripped as well due to the warm coolent from the SHMS circuling through the loop. After ~3h the ozone levels dropped below 0.3 which allowed the ARM to escort us in (https://logbooks.jlab.org/entry/3920045). The ARM kindly agreed to allow us to stay and revive as many systems as possible. We ended up bypassing the UPS and powering the cryo controllers and the gaussmeters directly from the power plugs avaiable in the SHMS hut (https://logbooks.jlab.org/entry/3920077). Heidi and Steve were in contact with Ellen and helped us recover both HMS and SHMS spectrometers. The techs were on hand to help with the vacuum that had increased in the spectrometers due to the incident. Ellen again made an access to troubleshoot the HMS Q1 and Q2 magnets (https://logbooks.jlab.org/entry/3920351). Many thanks to everyone for pitching in and getting us back up in such a short time.
* Yesterday Dave M made an escorted access to address the ice ball that was forming above the target chamber (https://logbooks.jlab.org/entry/3919519). This was initially delayed due to the ozone monitor but luckly it as during working hours when the Industrial Hygenist could come in and make a measurement with a hand held device. By the evening the ice ball had completely disappeared (https://logbooks.jlab.org/entry/3919858).
* collect production data until further notice (we have collected ~53% of the data at this kinematic point so far).
Unless something urgent comes up we will plan to meet again Monday at 10AM for the RC Daily Meeting.
This is a plain text email for clients that cannot display HTML. The full logentry can be found online at https://logbooks.jlab.org/entry/3920386
-------------- next part --------------
An HTML attachment was scrubbed...
More information about the Hallc_running